2.1 Class 8 Maths Question 3.
Solve the equation: 6 = z + 2

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Solution: We have 6 = z + 2 ⇒ 6 – 2 = z + 2 – 2 (subtracting 2 from each side) ⇒ 4 = z Thus, z = 4 is the required solution.